OpenStudy (anonymous):

Find the y – intercept of y – 7 = -8x A: 8/7 B: -7 C: 7 D: -8/7

jimthompson5910 (jim_thompson5910):

Plug in x = 0 and solve for y

jimthompson5910 (jim_thompson5910):

y – 7 = -8x y – 7 = -8(0) y - 7 = 0 I'll let you finish